OpenStudy (anonymous):

can someone explain steps to calculate the following: 9 (2/3)^1 times (1/3)^9 ===

OpenStudy (unklerhaukus):

\[9 \times{\left(\frac23\right)}^1 \times \left(\frac 13\right)^9\] ?

OpenStudy (anonymous):

how do you perform this math? can you explain?

OpenStudy (unklerhaukus):

so i have the right equation?

OpenStudy (unklerhaukus):

use these \[x^1=x\] \[\left(\frac ab\right)^n=\frac{a^n}{b^n}\]

OpenStudy (anonymous):

Supposely the result is .0004 for the problem ... why?

OpenStudy (unklerhaukus):

how did you get that

OpenStudy (anonymous):

My instructor That is why I am trying to figure out and can't

OpenStudy (unklerhaukus):

what number are you calculating?

OpenStudy (anonymous):

9 times (2/3)^1 times (1/3)^9 = .0004

OpenStudy (unklerhaukus):

\[9 \times{\left(\frac23\right)}^1 \times \left(\frac 13\right)^9\]\[=9\times\frac 23\times\frac 1{3^9}\]\[=\]

OpenStudy (unklerhaukus):

\[\neq0.0004\]

OpenStudy (anonymous):

I tried to do this manually right... I come up to (9/1) (2/3) (1/19683) = 18/59049 then is totally different result, right?

OpenStudy (unklerhaukus):

that is what i am too, \[=9\times\frac 23\times\frac 1{3^9}=\frac {18}{3^{10}}\approx0.0003048\]

OpenStudy (unklerhaukus):

are we solving the right equation?

OpenStudy (unklerhaukus):

i am getting*

OpenStudy (unklerhaukus):

\[(9+\frac23)\times\frac1{3^9}\approx0.0004911\]

OpenStudy (anonymous):

so that is where is coming from, i see now

OpenStudy (unklerhaukus):

but 0.0004911=0.0005 to one significant figure

OpenStudy (anonymous):

the part drives me nuts is that there is a difference when you round it, right?

OpenStudy (anonymous):

okay so I am allow to use a calculator that does not store or have memory, riight for the exam... I have a Casio FX-115s that will be allow right ... do you have any recommendations

OpenStudy (unklerhaukus):

ah, i have a Casio fx-82MS very similar , should be allowed in to exam

OpenStudy (unklerhaukus):

im not sure if your instructor made a rounding error or if the equation is different

OpenStudy (anonymous):

I believe he made a mistake... it is a few other mistakes he had done